please help me with this question. (-1+iroot3)(-1+iroot3)=: please help me with this question. (-1+iroot3)(-1+iroot3)=
Answer by Fombitz(1755)

(-1+sqrt(3)i)(-1+sqrt(3)i)
We use the FOIL method,
First : (-1)(-1)=1
Outer : (-1)(sqrt(3)i))=-sqrt(3)i
Inner : (sqrt(3)i)(-1))=-sqrt(3)i
Last : (sqrt(3)i)(sqrt(3)i)=3*-1=-3
Now sum all of the terms,
(-1+sqrt(3)i)(-1+sqrt(3)i)=1-sqrt(3)i-sqrt(3)i-3
(-1+sqrt(3)i)(-1+sqrt(3)i)=-2-2sqrt(3)i